Instructions





Upload photos of your answer sheet showing your name and solution. (2 items x 10 points)





1. Find A-B using the following matrices:





5





6 1





2





9





3





A=-3





5





8





B=|





81





4










3





3





4





67





-5





-1





2. Find AB using the following matrices:





A=





1 2





3





B=





+ Prepare answer





7





9





8





10





4 5 6





11





12






1
Expert's answer
2022-06-13T15:00:47-0400

1.


"A-B=\begin{pmatrix}
 5 & 6 & 1 \\
 -3 & 5 & 8 \\
 4 & 6 & 7 \\
\end{pmatrix}-\begin{pmatrix}
 2 & 9 &3 \\
 4 & 0 & 3 \\
 -5 & -1 & 3 \\
\end{pmatrix}"

"=\begin{pmatrix}
 5-2 & 6-9 & 1-3 \\
 -3-4 & 5-0 & 8-3 \\
 4-(-5) & 6-(-1) & 7-3 \\
\end{pmatrix}"

"=\begin{pmatrix}
 3 & -3 & -2 \\
 -7 & 5 & 5 \\
 9 & 7 & 4 \\
\end{pmatrix}"

2.

"AB=\begin{pmatrix}
 1 & 2 & 3 \\
 4 & 5 & 6 \\
\end{pmatrix}\begin{pmatrix}
 7 & 8 \\
 9 & 10 \\
 11 & 12 \\
\end{pmatrix}"

"=\begin{pmatrix}
 1(7)+2(9)+3(11) & 1(8)+2(10)+3(12) \\
 4(7)+5(9)+6(11) & 4(8)+5(10)+6(12) \\
 
\end{pmatrix}"


"=\begin{pmatrix}
 58 & 64 \\
 139 & 154
\end{pmatrix}"